Home‎ > ‎Crew 147‎ > ‎

117-AstronomyReport

Astronomy Report
01/17/2015
Scott MacPhee
Renee Garifi

Sky Conditions:
Clear
Wind Conditions:
Slight
Observation Start Time:
18:00
Observation End Time:
00:00
Summary:
Yesterday's observations of KC03C03248 were met with mixed results, as 
the dome ran out of power while tracking with the telescope. Prior to 
that, the autoguider was functioning well and the telescope remained on 
target with very little drift in images. Images obtained during the 
previous 2 evenings have been reduced to account for noise introduced in 
bias, dark, and flat frames.
For Jan 15th Approximately 40 exposures were collected, however the 70 
second exposures were insufficient to collect useful data on the target. 
Baseline analysis of potential comparison stars may be possible.
On Jan 16th Approximately 150 exposures were taken, however only 50 
images with 90 second exposure time are usable for analysis, as the dome 
did not rotate with the telescope.

As there are no transit candidates in the KELT database for this 
evening, today's Astronomy objectives involve performing photometric 
analysis on the data collected over the previous two nights and taking 
some astrophotos in the observatory later on. This will more readily 
allow us to process and graph transit light curves when one is observed 
in full.

Objects Viewed:
Lovejoy Comet should still be visible this evening. Imaging will be lead 
by Renee.

Problems Encountered:
Dome power loss. Midday test confirmed solar panels are recharging the 
dome.
Memory used by AstroImageJ (640MB) was insufficient to reduce images, 
thus was increased to 1000 MB. Software functioning as expected.

Starry Skies,
Renee & Scott